# A recommender-network perspective on the informational value of critics and crowds

### v1

@@ Abstract
How do the ratings of critics and amateurs compare and how can they be combined? Previous research has produced mixed results on the first question, while the second remains unanswered. We have created a new, unique dataset, with wine ratings from critics and amateurs, and simulated a recommender system using the weighted $k$-nearest-neighbor algorithm. We then formalized the advice-seeking network spanned by that algorithm (i.e., who advises whom?) and studied people’s relative influence. We find that critics are more consistent than amateurs, and thus their advice is more predictive than advice from amateurs. Getting advice from both groups can further boost performance. Our network-theoretic approach allows us to identify influential critics, talented amateurs, and the information flow between groups. Our results provide evidence about the informational function of critics, while our framework is broadly applicable and can be leveraged to devise good decision strategies and more transparent recommender systems.
@@ 1 Introduction
Whether it is a film currently in cinemas, a restaurant that just opened, or vintage wine, people like to voice their judgments on matters of taste. Even more so, for a few select individuals—critics—expressing judgments on matters of taste has turned from a fun past time to a profession. Critics are employed in the daily and weekly press to assess restaurants, theater, movies, and wine labels, while some even run popular television shows. It has been argued that critics function as information producers, generating information about products that becomes available to the wider public [43, 71]. This key function of critics is especially pronounced on websites such as Metacritic and Rotten Tomatoes, which have based their business model entirely on devising and displaying scores that summarize critics’ opinions, or when they are hired as judges in prestigious competitions. In fact, it has been shown that the judgments of influential critics can predict [29] or even alter the overall financial performance of products [3, 12]. But are the evaluations of critics more informative than those of amateur consumers, and if so, why?
Even though there is clear evidence for the sizable market impact of at least some critics, the informational value of their opinions for the broader public has often been contested. Popular lore and scholars in different traditions have argued that in matters of taste, there is often a big, unresolvable divide between critics and the general public [30]. Some would suggest that critics and other experts engage in hairsplitting over nuances that are often lost on the rest of the world, who consider expert opinions high-brow, and ivory tower, if not altogether imaginary or downright boring [34, 14]. Others emphasize that general audiences lack the ability to assess sophisticated (cultural) products, and only people who have trained their senses are capable of doing so [2]. Regardless of the point of view adopted, other people may rarely care about what critics think, whatever theirs standards of taste may be. Thus, although critics’s opinions may be impactful (e.g. because they bring attention to certain categories of cultural or experience products [72]), they could also be less informative than the opinions of representative individuals in the crowd or even a randomly selected person from the street [41]. That is, the opinions of a critic or a group of critics might be less correlated with and, most importantly, less predictive of the tastes of people in the wider public, than those of other amateurs.
@@
Several questions about the ways in which critics can inform and influence the broader public remain unanswered. First, are their opinions more informative than those of amateurs, and if so, why? Second, is it possible to identify informative critics and talented amateurs who have the potential to become critics? Third, how can the opinions of critics and amateurs be best combined and how would information flow between the two groups in an efficient advice network? To answer these questions we need a methodology that allows us to measure the informational value of the opinions of different groups of individuals for others as well as the individuals’ potential to influence others.
@@
Since the advent of the internet and social media, the informational landscape in the wine world has been changing swiftly. Vivino has emerged as a reliable alternative to critics for getting access to information about the quality of different wines, and as with other opinion aggregation websites (e.g. IMDB, Yelp, Amazon) it gives wine consumers a way to consult a single information source to inform their choices, bringing forth a democracy of taste where each individual in the user base can contribute to the aggregate wine ratings [46]. Personalization algorithms can search for similarity patterns between one person and others in rating databases, and they can receive recommendations based on a number of apparently similar people from across the globe, thus circumventing critics altogether [1, 67]. To examine how the opinions of critics relate to those of amateurs we created a new, unique dataset consisting of the ratings from both renowned wine critics and regular wine consumers (i.e., amateurs, non-professionals). We obtained critics’ data from Bordoverview, a website summarizing the en primeur ratings (first sampling of a production year) on Bordeaux wines, and matched these ratings with amateur data on the same wines from Vivino. The resulting dataset has the properties needed to assess how wine critics and amateurs can inform and influence others and how their judgements on matters of taste can be best combined.
@@ 4.6 Conclusion
In vino veritas—in wine, there is truth—says an old Latin adage. We found that critics’ judgements are indeed valuable in helping amateurs identify good wines, more so than the opinion of most other amateurs. Still, there is scope for combining the opinions of both critics and amateurs, and for identifying the most influential critics and talented amateurs, whose tastes appear to be informative for many other individuals. Going beyond wine, the methods we developed are modular and generic and can be readily applied in any dataset where different groups of people have rated a number of items, even when there are discrepancies in the number of items evaluated, to tackle long-lasting research questions in the social and management sciences. Further, they can be also used as a tool for improving people’s understanding of key collaborative filtering algorithms [77] by enhancing the transparency and interpretability of the recommendation process [35], and can help people hone in on the right decision strategies when seeking a good item in matters of taste.

### v2

@@ Abstract
How do the ratings of critics and amateurs compare and how can they be combined? Previous research has produced mixed results on the first question, while the second remains unanswered. We have created a new dataset with wine ratings from critics and amateurs, and simulated a recommender system using the weighted $k$-nearest-neighbor algorithm. We then formalized the advice-seeking network spanned by that algorithm (i.e., who advises whom?) and studied people’s relative influence. We find that critics are more consistent than amateurs, and thus their advice is more predictive of people’s taste than advice from amateurs. Getting advice from both groups can further boost performance, but only by a small margin. Our network-theoretic approach allows us to identify influential critics, talented amateurs, and the information flow between groups. Our results provide evidence about the informational function of critics, while our framework is broadly applicable and can be leveraged to devise good decision strategies and more transparent recommender systems.
@@ 1 Introduction
Whether it is a film currently in cinemas, a restaurant that just opened, or vintage wine, people like to voice their judgments on matters of taste. Even more so, for a few select individuals—critics—expressing judgments on matters of taste has turned from a fun past time to a profession. Critics are employed in the daily and weekly press to assess restaurants, theater, movies, and wine labels, while some even run popular television shows. Critics may function as information producers, supplying readily accessible information about products to the wider public [43, 71]. This key function of critics is especially pronounced on websites such as Metacritic and Rotten Tomatoes, which have based their business model on collecting critics’ opinions and devising scores that summarize their opinions, or when they are hired as judges in prestigious competitions. So far, it has been shown that the judgments of influential critics can predict [29] or even alter the overall financial performance of products [3, 12]. But are the evaluations of critics more informative than those of amateur consumers, and if so, why?
Even though there is clear evidence for the sizable market impact of at least some critics, the informational value of their opinions for the broader public has often been contested. Popular lore and scholars in different traditions have argued that in matters of taste, there is often a big, unresolvable divide between critics and the general public [30]. Some would suggest that critics and other experts engage in hairsplitting over nuances that are often lost on the rest of the world, who consider expert opinions high-brow, and ivory tower, if not altogether imaginary or downright boring [34, 14]. Others emphasize that general audiences lack the ability to assess sophisticated (cultural) products, and only people who have trained their senses are capable of doing so [2]. Regardless of the point of view adopted, other people may rarely care about what critics think, whatever theirs standards of taste may be. Thus, although critics’s opinions may be impactful (e.g. because they bring attention to certain categories of cultural or experience products [72]), they could also be less informative than crowd-sourced aggregate scores, the opinions of representative individuals in the crowd or even a randomly selected person from the street [41]. That is, the opinions of a critic or a group of critics might be less correlated with and, most importantly, less predictive of the tastes of people in the wider public, than those of other amateurs.
@@
Several questions about the ways in which critics and/or amateur crowds can inform and influence the broader public remain unanswered. First, are the opinions of critics more informative than those of amateurs/amateur crowds, and if so, why? Second, is it possible to identify the most informative critics and talented amateurs who have the potential to become critics? Third, how can the opinions of critics and amateur crowds be best combined and how would information flow between the two groups in an efficient advice network? To answer these questions we need a methodology that would allow us to measure the informational value of the opinions of different individuals or groups for others as well as the their potential to influence them.
@@
Since the advent of the internet and social media, the informational landscape in the wine world has been changing swiftly. Vivino has emerged as a reliable alternative to critics for getting access to information about the quality of different wines, and as with other opinion aggregation websites (e.g. IMDB, Yelp, Amazon) it gives wine consumers a way to consult a single information source to inform their choices, bringing forth a democracy of taste where each individual in the user base can contribute to the aggregate wine ratings [46]. Personalization algorithms can search for similarity patterns between one person and others in rating databases, and they can receive recommendations based on a number of apparently similar people from across the globe, thus circumventing critics altogether [1, 67]. To examine how the opinions of critics relate to those of amateurs we created a new dataset consisting of the ratings from both renowned wine critics and regular wine consumers (i.e., amateurs, non-professionals). We obtained critics’ data from Bordoverview, a website summarizing the en primeur ratings (first sampling of a production year) on Bordeaux wines, and matched these ratings with amateur data on the same wines from Vivino. The resulting dataset has the properties needed to assess how wine critics and amateurs can inform and influence others and how their judgements on matters of taste can be best combined.
@@ 4.6 Conclusion
In vino veritas—in wine, there is truth—says an old Latin adage. We found that critics’ judgements are indeed valuable in helping amateurs identify good wines, more so than the opinion of most other amateurs. Still, there is scope for combining the opinions of both critics and amateurs, and for identifying the most influential critics and talented amateurs, whose tastes appear to be informative for many other individuals. Going beyond wine, the methods we developed are modular and generic and can be readily applied in any dataset where different groups of people have rated a number of items, even when there are discrepancies in the number of items evaluated, to tackle long-lasting research questions in the social, behavioral and management sciences. Further, they can be also used as a tool for improving people’s understanding of key collaborative filtering algorithms [77] by enhancing the transparency and interpretability of the recommendation process [35], and can help people hone in on the right decision strategies when seeking a good item in matters of taste.